OpenStudy (anonymous):

Six is the geometric mean of 18 and what number? 2 3 4 9

OpenStudy (anonymous):

2

OpenStudy (anonymous):

say 6 is geometric mean of 18 and x => 6^2 = (18)(x) => x = 36/18 = 2

OpenStudy (mertsj):